OpenStudy (anonymous):

a cube 10 cm on edge can hold a liter of water. A foot is about 30 cm. How many liters of water can a cube 30 cm on edge hold?

OpenStudy (richyw):

\[(10\text{cm})^3=1000\text{cm}^3=1\text{L}\]\[(30)^3=(3\times 10)^3=(3^3\times 10^3)=9\text{L}\]